OpenStudy (anonymous):

Carlos and Kathryn are selling pies for a school fundraiser. Customers can buy apple pies and lemon meringue pies. Carlos sold 14 apple pies and 6 lemon meringue pies for a total of $266. Kathryn sold 6 apple pies and 10 lemon meringue pies for a total of $218. Find the cost each of one apple pie and one lemon meringue pie.

OpenStudy (anonymous):

14x+6y=266 and 6x+10y=218

OpenStudy (anonymous):

7x+3y=133 and 2x+5y=109 right?

OpenStudy (anonymous):

I'm not sure what to do next though.

OpenStudy (phi):

those look good. I would simplify both of them by divide both sides by 2

OpenStudy (phi):

7x+3y=133 2x+5y=109 if you make the numbers in front of the y (for example) the same , we can solve I would multiply the top equation (both sides) by 5 and the bottom by 3 what do you get for the two new equations?

OpenStudy (phi):

notice I wanted the greatest common factor (like finding a "common denominator") of 3 and 5. i.e. 15

OpenStudy (anonymous):

35x+15y and 6x+15y?

OpenStudy (phi):

except you have to write the *entire* equation. (both sides)

OpenStudy (anonymous):

but is 35x+15y right?

OpenStudy (phi):

yes, you did the left half of each equation correctly

OpenStudy (anonymous):

okay so you multiply 133 by 5 and 109 by 3?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok so then we have 35x+15y=665 and 6x+15y=327

OpenStudy (phi):

35x+15y=665 6x+15y=327 now subtract the bottom equation from the top equation or if you like, multiply the bottom equation by -1 (both sides), then add the equations

OpenStudy (phi):

I notice a small mistake you started with 6x+10y=218 when you simplify by dividing both sides by 2 you should get 3x + 5y = 109 (you wrote down 2x ).

OpenStudy (anonymous):

I accidentally wrote 2 instead of 3 haha

OpenStudy (anonymous):

I noticed that too

OpenStudy (anonymous):

I'm kinda stuck on these last equations though.

OpenStudy (phi):

that changes the equations to 35x+15y=665 9x+15y=327

OpenStudy (anonymous):

35-9=26 and then the 15 cancels out and 665-327=338

OpenStudy (phi):

yes, that is the correct idea, but leave in the letter you are doing 35x -9x + 15y - 15y = 665-327

OpenStudy (phi):

35 x's take away 9 x's leaves 26 x's i.e. 35x -9x + 15y - 15y = 665-327 simplifies to 26x = 338

OpenStudy (phi):

now divide both sides by 26

OpenStudy (anonymous):

like 338 divided by 26? that's 13

OpenStudy (phi):

yes. you are doing \[ \frac{26x}{26} = \frac{338}{26} \] now simplify both sides

OpenStudy (phi):

on the left side 26/26 is 1, and 1*x is just x on the right side you get 13 so after simplifying you get x= 13

OpenStudy (anonymous):

perfect and then y=14 :) thank you

OpenStudy (anonymous):

I have 2 other questions just like this, could you help me again? I'm starting to get the hang of it haha

OpenStudy (phi):

ok, but please make a new post.

OpenStudy (anonymous):

gotcha :)